You are here: TWiki > SoftwareTeam Web>Meetings>2011-10-17 (17 Oct 2011, AlainRoy?)

OSG Software Team Meeting for 17-Oct-2011

Attending

  • Cartwright, Tim
  • Chism, Elizabeth
  • Dost, Jeff
  • Gore, Brook
  • Kronenfeld, Scot
  • Roy, Alain
  • Mambelli, Marco
  • Selmeci, Mat
  • Sim , Alex
  • Sfiligoi, Igor
  • Strain, Doug
  • Thapa, Suchandra
  • Weichel, Jim

Apologies

State of documentation

All documentation tasks are now tickets in JIRA. Each one has a component titled "documentation". During the meeting, we went through the open tickets and made sure everyone knows what they have. If you run across documentation needs that aren't a ticket, please create a ticket or send an email to Jim about it.

Make sure you update the document on what pages have been moved to the new documentation.

Jim will make a ticket to sort out documentation that is in the storage twiki. The overall criteria is: 1) If it's release specific, it moves to the Documentation.Release3 twiki. If it's general-purpose documentation, it moves to the Documentation twiki. If it's specific to the old storage subgroup (like old meeting minutes), it stays there. In the future, no new documentation will go into the storage twiki.

State of the CE

settle environment of job (from mailing list discussion)

General agreement.

Status Reports

  • Alex: Changing a lot of things in Bestman. Dropping dependent libraries that aren't needed. Changing a lot of internal stuff to deal with that. This is time consuming, but in progress. Dropping old SAML version. Alex said that this new version will only work with Java 7, but Alain said that it's unworkable: it has to work with Java 6 and 7. While we expect we'll be using Java 7 in the future, we must continue to work with Java 6 for quite a while. Alex said he'll look into it.
  • Doug: Primarily focused on getting new storage documentation into order.
  • Igor: Terrence found some problems when running on a larger Condor pool. He needs to report them soon, but hasn't yet.
  • Jeff: GridFTP HDFS RPM is complete. Brian wants to push some GridFTP init script changes upstream to Globus, and that's in progress. There's a problem with upgrading. Will be working on process for upgrading from CMS HDFS to OSG repo.
  • Marco: Trying to reproduce errors that Scot had with cacet-expiry RSV probe. Now has fermicloud access.
  • Mat: Spent most of last week working on the LIGO packages for Scientific Linux 6 and clearing out old tickets: Koji documentation and Koji garbage collection.
  • Scot: Working on RSV documentation. Have a couple of small RSV enchancements in progress.
  • Suchandra: Test install of CE and RSV last week. Listening to feedback from Tony to improve osg-configure. Ran into errors in RSV probes for SRM. Will push out new osg-configure release today.
  • Tim: Worked on LIGO packages for Debian 6 last week, and they're probably done. Will resume work on our test framework

Other business

None

Topic revision: r1 - 17 Oct 2011 - 21:32:11 - AlainRoy?
Hello, TWikiGuest
Register

 
TWIKI.NET

TWiki | Report Bugs | Privacy Policy

This site is powered by the TWiki collaboration platformCopyright by the contributing authors. All material on this collaboration platform is the property of the contributing authors..